# Rethinking Disentanglement under Dependent Factors of Variation

### v2

@@ 4 Measuring Disentanglement via Minimality and Sufficiency
Given a representation $z$ of an input $x$ and a target variable $y$, we say that $z$ is minimal with respect to $y$ if it satisfies $I(z;x|y)=0$ or, equivalently, if $I(z;y)=I(z;x)$, i.e., $z$ contains information only about $y$.
@@
Given a representation $z$ of an input $x$ and a target variable $y$, we say that $z$ is sufficient with respect to $y$ if it satisfies $I(y;x|z)=0$ or, equivalently, if $I(z;x)=I(y;x)$, i.e., $z$ contains all information about $y$.
@@ 4.1 Connection between Disentanglement and Minimality and Sufficiency
Let $\bm{y}=\{y_{l}\}_{l=1}^{n}$ denote a set of factors and $\bm{z}=\{z_{j}\}_{j=1}^{m}$ a representation. Then, $z_{j}$ is a minimal representation of $y_{i}$ if and only if $z_{j}$ is factors-invariant with respect to $y_{i}$ and nuisances-invariant. Equivalently, we have that:  [^3] Note that $\Longleftrightarrow$ denotes ”if and only if”, whereas $\leftrightarrow$ represents a dependency structure in a Markov chain.
$(\bm{x}\leftrightarrow y_{i}\leftrightarrow z_{j})\Longleftrightarrow(\bm{\tilde{y}}_{i}\leftrightarrow y_{i}\leftrightarrow z_{j})\,\land\,(\bm{n}\leftrightarrow y_{i}\leftrightarrow z_{j})$
where $\bm{\tilde{y}}_{i}=\{y_{l}\}_{l\neq i}$. Therefore, satisfying minimality is equivalent to jointly satisfying factors-invariance and nuisances-invariance. Intuitively, if $z_{j}$ is fully determined by $y_{i}$ (i.e., minimal), then neither the remaining factors $\bm{\tilde{y}}_{i}$ nor the nuisances $\bm{n}$ influence $z_{j}$ once $y_{i}$ is known—captured by factors-invariance and nuisances-invariance, respectively. Conversely, since the input $\bm{x}$ is completely defined by $\bm{y}$ and $\bm{n}$, the implication also holds in the reverse direction. A formal proof of this theorem is provided in Appendix A.
@@
Let $y_{i}$ denote a factor and $\bm{z}=\{z_{l}\}_{l=1}^{m}$ a representation. Then, $z_{j}$ is a sufficient representation of $y_{i}$ if and only if $z_{j}$ is representations-invariant for $y_{i}$ and $\bm{z}$ is explicit for $y_{i}$. Equivalently, we have that:
where $\bm{\tilde{z}}_{j}=\{z_{l}\}_{l\neq j}$. Thus, satisfying sufficiency is equivalent to jointly satisfying representations-invariance and explicitness. Intuitively, if $y_{i}$ is fully determined by $z_{j}$ independently of $\bm{x}$ (i.e., sufficiency), then there is no information about $y_{i}$ in $\bm{z}$ that is not already contained in $z_{j}$ (representations-invariance), and since $z_{j}\in\bm{z}$, this also implies that $y_{i}$ is fully captured by $\bm{z}$ (explicitness). Conversely, if all the information in $\bm{z}$ relevant to $y_{i}$ is already present in $z_{j}$ (representations-invariance), and $\bm{z}$ captures all information about $y_{i}$ (explicitness), then $z_{j}$ must fully describe $y_{i}$—that is, $z_{j}$ is sufficient. A formal proof of this theorem is provided in Appendix A.
@@ 4.2 Why Measuring Disentanglement via Minimality and Sufficiency?
  • Minimality vs. Factors-Invariance + Nuisances-Invariance: When evaluating whether a representation is disentangled, the true objective is to determine whether each of its variables is influenced by a single factor—regardless of the presence of other factors or nuisances. It is not sufficient for a variable to be aligned with a specific factor if it is also significantly affected by nuisance variables. Similarly, a variable that is invariant to nuisances but still influenced by multiple factors cannot be considered disentangled. Therefore, we argue that these two properties should be assessed jointly, and—as established in Theorem 1—this can be effectively achieved by measuring minimality.
@@ 4.3 Defining metrics for Minimality and Sufficiency
At the beginning of this section, we defined when a representation is minimal or sufficient. In practice, however, representations are rarely perfectly minimal or sufficient. It is therefore desirable to design continuous metrics that quantify how close a representation is to these ideals. Moreover, while minimality and sufficiency can be defined for each pair $(z_{j},y_{i})$, it is often useful to summarize these values into a single score that characterizes the entire representation $\bm{z}$. We next introduce metrics that meet these requirements.
@@
Recall that a variable $z_{j}$ is minimal with respect to a factor $y_{i}$ if $I(z_{j};\bm{x}\mid y_{i})=I(z_{j};x)-I(z_{j};y_{i})=0$. To improve interpretability and facilitate comparison across metrics, we normalize this value by $I(z_{j};\bm{x})$ to ensure it lies in $[0,1]$. We thus define the minimality of $z_{j}$ with respect to $y_{i}$ as
$m_{ij}=1-\frac{I(z_{j};\bm{x}\mid y_{i})}{I(z_{j};\bm{x})}=\frac{I(z_{j};y_{i})}{I(z_{j};x)}.$ (5)
Here, $m_{ij}$ reflects the fraction of information in $z_{j}$ that is attributable to factor $y_{i}$. While these values provide factor-wise insights, it is common to report a single scalar as a disentanglement score. Assuming, as in Section 3, a one-to-one correspondence between $z_{j}$ and a specific $y_{i}$, we define $\hat{m}_{j}=\max_{i}m_{ij}$, measuring the minimality of $z_{j}$ relative to its most associated factor. The overall minimality score is then obtained by averaging across all representation variables: $\bar{m}=\frac{1}{m}\sum_{j}\hat{m}_{j}$.
@@
Analogously, a variable $z_{j}$ is sufficient for a factor $y_{i}$ if $I(y_{i};\bm{x}\mid z_{j})=I(y_{i};\bm{x})-I(y_{i};z_{j})=0$. Applying the same normalization yields the sufficiency of $z_{j}$ with respect to $y_{i}$:
$s_{ij}=1-\frac{I(y_{i};\bm{x}\mid z_{j})}{I(y_{i};\bm{x})}=\frac{I(y_{i};z_{j})}{I(y_{i};x)}.$ (6)
Thus, $s_{ij}$ measures the proportion of information about $y_{i}$ captured by $z_{j}$. Following the same reasoning as for minimality, we compute $\hat{s}_{i}=\max_{j}s_{ij}$, which evaluates how well the most informative representation variable predicts $y_{i}$. Averaging across factors yields the overall sufficiency score: $\bar{s}=\frac{1}{n}\sum_{i}\hat{s}_{i}$.
4.4 Calculating Minimality and Sufficiency in Practice
Calculating the mutual information terms from the previous section can be intractable, particularly when the factor $y_{i}$ or the representation $z_{j}$ is continuous. To address this, we proceed as follows: (i) we discretize $z_{j}$ by binning, a common practice in representation learning (Shwartz-Ziv & Tishby, 2017; Locatello et al., 2019b; Carbonneau et al., 2022); and (ii) in the less common case where $y_{i}$is not categorical, we also discretize it by binning, while leaving it unchanged when it is categorical. Once both $y_{i}$ and $z_{j}$ are categorical, we obtain: (i) $I(z_{j};y_{i})$can be computed directly; (ii) $I(z_{j};x)=H(z_{j})$, since $z_{j}$ is fully defined by $x$ and therefore $H(z_{j}|x)=0$; and (iii) $I(y_{i};x)=H(y_{i})$, by the same reasoning. In all experiments presented in the following sections, we discretize the representations using 15 bins, while the factors are categorical in every case.
@@ 5.1 Disentanglement under Dependence between Factors of Variation
for $i=1,\dots,n$, where $\delta\in\left[\frac{1}{n},1\right]$ is a hyperparameter that controls the level of independence among the factors, and the function $\mathcal{B}:\mathbb{R}\to\left\{0,1,\dots,K\right\}$ bins the continuous values into $K$ discrete classes. In particular: (i) when $\delta=1$, each $y_{i}$ depends solely on $\epsilon_{i}$, resulting in independent factors; and (ii) when $\delta=\frac{1}{n}$, each $y_{i}$ is equally influenced by all elements of $\bm{\epsilon}$, yielding maximal dependence.
@@
Figure 2 shows the results of various disentanglement metrics for $n=4$ and $K=5$ across different values of $\alpha$ and $\delta$. We observe that Minimality and Sufficiency are the only metrics that consistently satisfy the desirable properties outlined above. More specifically, we can see that: (i) Metrics such as MIG, SAP, Disentanglement, and Completeness fail to reach a value of one when $\alpha=1$ and $\delta=1$, i.e., even in the fully disentangled and independent case; (ii) Metrics such as FactorVAE, MIG, SAP, and Modularity never attain their maximum values as $\delta=\frac{1}{n}$, meaning they fail to correctly reflect disentanglement when correlations are present; and (iii) Disentanglementand Completeness remain largely invariant with respect to $\delta$, contrary to the expected behavior discussed earlier.
@@
Figure 2: Comparison of different metrics (y-axis) across varying values of $\alpha$ (x-axis) and $\delta$ (color). Colors range from dark blue (highest dependence, $\delta=\frac{1}{n}$) to dark red (lowest independence, $\delta=1$).
@@ 6 Neural Representations Experiments
To this end, we train models known to induce disentangled representations on three widely used datasets: DSprites (Matthey et al., 2017), MPI3D (Gondal et al., 2019), and Shapes3D (Burgess & Kim, 2018). For each dataset, we explore six encoder architectures (detailed in Appendix C) combined with six disentanglement losses: $\beta$-VAE (Higgins et al., 2017), AnnealedVAE (Burgess et al., 2018), $\beta$-TCVAE (Chen et al., 2018), FactorVAE (Kim & Mnih, 2018), Ada-GVAE (Locatello et al., 2020), and HFSVAE (Roth et al., 2022), yielding 36 encoders per dataset. All models are trained using the hyperparameters listed in Appendix B.
@@ 6.2 Interventional Robustness
Another motivation for disentangled representations is their robustness to interventions on the underlying factors. Specifically, if a factor $y_{i}$ is modified while all others remain fixed, the resulting change in the representation should be confined to its corresponding dimension $z_{j}$. This property is particularly critical for controlled generative models: if it holds, modifying $z_{j}$ ensures that the generated data changes exclusively in the associated factor $y_{i}$. To quantify this, Suter et al. (2019) introduced the Interventional Robustness Score (IRS), which quantifies the robustness of a representation under changes in the factors. We next compute the rank correlation between IRS and all considered metrics.
@@
Figure 5: Rank correlation between disentanglement metrics and Interventional Robustness Score across different levels of factor dependence.
@@ Acknowledgments
This work has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101007666, MCIN/AEI/10.13039/501100011033 under Grant PID2021-126061OB-C44, and the Government of Aragón (Grant Group T36 23R).

### v3

@@ 4 Measuring Disentanglement via Minimality and Sufficiency
Given a representation variable $Z$ of an input $X$ and a target variable $Y$, we say that $Z$ is minimal with respect to $Y$ if it satisfies $I(Z;X|Y)=0$ or, equivalently, if $I(Z;Y)=I(Z;X)$, i.e., $Z$ contains information only about $Y$.
@@
Given a representation variable $Z$ of an input $X$ and a target variable $Y$, we say that $Z$ is sufficient with respect to $Y$ if it satisfies $I(Y;X|Z)=0$ or, equivalently, if $I(Z;X)=I(Y;X)$, i.e., $Z$ contains all information about $Y$.
@@ 4.1 Connection between Disentanglement and Minimality and Sufficiency
Let $Y=\{Y_{k}\}_{k=1}^{n}$ denote a set of factors and $Z=\{Z_{j}\}_{j=1}^{m}$ a representation. If $Z_{j}$ is a minimal representation of $Y_{i}$, it follows that $Z_{j}$ is factors-invariant with respect to $Y_{i}$ and nuisances-invariant. Equivalently, we have that:
$(X\leftrightarrow Y_{i}\leftrightarrow Z_{j})\Longrightarrow(\tilde{Y}_{i}\leftrightarrow Y_{i}\leftrightarrow Z_{j})\,\land\,(N\leftrightarrow Y_{i}\leftrightarrow Z_{j})$
where $\tilde{Y}_{i}=\{Y_{k}\}_{k\neq i}$. Therefore, satisfying minimality implies jointly satisfying factors-invariance and nuisances-invariance. Intuitively, if $Z_{j}$ is fully determined by $Y_{i}$ (i.e., minimal), then neither the remaining factors $\tilde{Y}_{i}$ nor the nuisances $N$ influence $Z_{j}$ once $Y_{i}$ is known—captured by factors-invariance and nuisances-invariance, respectively. A formal proof of this theorem is provided in Appendix A.
@@
Let $Y_{i}$ denote a factor and $Z=\{Z_{k}\}_{k=1}^{m}$ a representation. Then, $Z_{j}$ is a sufficient representation of $Y_{i}$ if and only if $Z_{j}$ is representations-invariant for $Y_{i}$ and $Z$ is explicit for $Y_{i}$. Equivalently, we have:  [^1] Note that $\Longleftrightarrow$ denotes ”if and only if”, whereas $\leftrightarrow$ represents a dependency structure in a Markov chain.
$(X\leftrightarrow Z_{j}\leftrightarrow Y_{i})\Longleftrightarrow(\tilde{Z}_{j}\leftrightarrow Z_{j}\leftrightarrow Y_{i})\,\land\,(X\leftrightarrow Z\leftrightarrow Y_{i})$
where $\tilde{Z}_{j}=\{Z_{k}\}_{k\neq j}$. Thus, satisfying sufficiency is equivalent to jointly satisfying representations-invariance and explicitness. Intuitively, if $Y_{i}$ is fully determined by $Z_{j}$ independently of $X$ (i.e., sufficiency), then there is no information about $Y_{i}$ in $Z$ that is not already contained in $Z_{j}$ (representations-invariance), and since $Z_{j}\in Z$, this also implies that $Y_{i}$ is fully captured by $Z$ (explicitness). Conversely, if all the information in $Z$ relevant to $Y_{i}$ is already present in $Z_{j}$ (representations-invariance), and $Z$ captures all information about $Y_{i}$ (explicitness), then $Z_{j}$ must fully describe $Y_{i}$—that is, $Z_{j}$ is sufficient. A formal proof of this theorem is provided in Appendix A.
@@ 4.2 Why Measuring Disentanglement via Minimality and Sufficiency?
  • Minimality vs. Factors-Invariance + Nuisances-Invariance: When evaluating whether a representation is disentangled, the true objective is to determine whether each of its variables is influenced by a single factor—regardless of the presence of other factors or nuisances. It is not sufficient for a variable to be aligned with a specific factor if it is also significantly affected by nuisance variables. Similarly, a variable that is invariant to nuisances but still influenced by multiple factors cannot be considered disentangled. Therefore, we argue that these two properties should be assessed jointly, and—as established in Theorem 1—this can be effectively achieved by measuring minimality. Moreover, analyzing factors and nuisances separately overlooks joint effects, where correlations between remaining factors and nuisances affect $z_{j}$ collectively despite having no independent influence. Minimality, by contrast, captures these underlying correlations. Furthermore, from a practical standpoint, labels for nuisance variables are seldom available, which prevents the explicit calculation of nuisance-invariance.
@@ 4.3 Defining metrics for Minimality and Sufficiency
At the beginning of this section, we defined when a representation is minimal or sufficient. In practice, however, representations are rarely perfectly minimal or sufficient. It is therefore desirable to design continuous metrics that quantify how close a representation is to these ideals. Moreover, while minimality and sufficiency can be defined for each pair $(Z_{j},Y_{i})$, it is often useful to summarize these values into a single score that characterizes the entire representation $Z$. We next introduce metrics that meet these requirements.
@@
Recall that a variable $Z_{j}$ is minimal with respect to a factor $Y_{i}$ if $I(Z_{j};X\mid Y_{i})=I(Z_{j};X)-I(Z_{j};Y_{i})=0$. To improve interpretability and facilitate comparison across metrics, we normalize this value by $I(Z_{j};X)$ to ensure it lies in $[0,1]$. We thus define the minimality of $Z_{j}$ with respect to $Y_{i}$ as
$m_{ij}=1-\frac{I(Z_{j};X\mid Y_{i})}{I(Z_{j};X)}=\frac{I(Z_{j};Y_{i})}{I(Z_{j};X)}.$ (5)
Here, $m_{ij}$ reflects the fraction of information in $Z_{j}$ that is attributable to factor $Y_{i}$. While these values provide factor-wise insights, it is common to report a single scalar as a disentanglement score. Assuming, as in Section 3, a one-to-one correspondence between $Z_{j}$ and a specific $Y_{i}$, we define $\hat{m}_{j}=\max_{i}m_{ij}$, measuring the minimality of $Z_{j}$ relative to its most associated factor. The overall minimality score is then obtained by averaging across all representation variables:
$\bar{m}=\frac{1}{m}\sum_{j}\hat{m}_{j}$ (6)
@@
Analogously, a variable $Z_{j}$ is sufficient for a factor $Y_{i}$ if $I(Y_{i};X\mid Z_{j})=I(Y_{i};X)-I(Y_{i};Z_{j})=0$. Applying the same normalization yields the sufficiency of $Z_{j}$ with respect to $Y_{i}$:
$s_{ij}=1-\frac{I(Y_{i};X\mid Z_{j})}{I(Y_{i};X)}=\frac{I(Y_{i};Z_{j})}{I(Y_{i};X)}.$ (7)
Thus, $s_{ij}$ measures the proportion of information about $Y_{i}$ captured by $Z_{j}$. Following the same reasoning as for minimality, we compute $\hat{s}_{i}=\max_{j}s_{ij}$, which evaluates how well the most informative representation variable predicts $Y_{i}$. Averaging across factors yields the overall sufficiency score:
$\bar{s}=\frac{1}{n}\sum_{i}\hat{s}_{i}$ (8)
4.4 Computing Minimality and Sufficiency in Practice
Calculating the mutual information terms requires access to the ground-truth factor labels $Y$. Furthermore, exact computation can be intractable, particularly when the factor $Y_{i}$ or the representation $Z_{j}$ is continuous. To address this, we proceed as follows: (i) we discretize $Z_{j}$ by binning, a common practice in representation learning (Shwartz-Ziv and Tishby, 2017; Locatello et al., 2019b; Carbonneau et al., 2022); and (ii) in the less common case where $Y_{i}$ is not categorical, we also discretize it by binning, while leaving it unchanged when it is categorical. Once both $Y_{i}$ and $Z_{j}$ are categorical, we obtain: (i) $I(Z_{j};Y_{i})$can be computed directly; (ii) $I(Z_{j};X)=H(Z_{j})$, since $Z_{j}$ is fully defined by $X$ and therefore $H(Z_{j}|X)=0$; and (iii) $I(Y_{i};X)=H(Y_{i})$, by the same reasoning.
@@ 5.1 Disentanglement under Dependence between Factors of Variation
for $i=1,\dots,n$, where $\delta\in\left[\frac{1}{n},1\right]$ is a hyperparameter that controls the level of independence among the factors, and the function $\mathcal{B}:\mathbb{R}\to\left\{1,2,\dots,K\right\}$ bins the continuous values into $K$ discrete classes. In particular: (i) when $\delta=1$, each $y_{i}$ depends solely on $\epsilon_{i}$, resulting in independent factors; and (ii) when $\delta=\frac{1}{n}$, each $y_{i}$ is equally influenced by all elements of $\bm{\epsilon}$, yielding maximal dependence.
@@
Figure 2 shows the results of various disentanglement metrics for $n=4$ and $K=5$ across different values of $\alpha$ and $\delta$. While we report results using a fixed discretization of 20 bins and 5000 data samples, we verify in Appendices B and C that the metrics’ performance is largely insensitive to changes in bin counts and number of samples, respectively. We observe that Minimality and Sufficiency are the only metrics that consistently satisfy the desirable properties outlined above. More specifically, we can see that: (i) Metrics such as MIG, SAP, Disentanglement, and Completeness fail to reach a value of one when $\alpha=1$ and $\delta=1$, i.e., even in the fully disentangled and independent case; (ii) Metrics such as FactorVAE, MIG, SAP, and Modularity never attain their maximum values as $\delta=\frac{1}{n}$, meaning they fail to correctly reflect disentanglement when correlations are present; and (iii) Disentanglementand Completeness remain largely invariant with respect to $\delta$, contrary to the expected behavior discussed earlier.
@@
Figure 2: Comparison of different metrics (y-axis) across varying values of $\alpha$ (x-axis) and $\delta$ (color). Colors range from dark blue (highest dependence, $\delta=\frac{1}{n}$) to dark red (lowest dependence, $\delta=1$).
@@ 5.2 Disentanglement under the Presence of Nuisances
Random text here
5.3 Scalability with Respect to Number of Factors
Figure 4: Time (in seconds) for calculating the different metrics (y-axis) for different number of factors of variation (x-axis).
To assess the practical feasibility of these metrics in high-dimensional scenarios, we evaluate their runtime performance as the number of generative factors increases (Figure 4). Regarding the experimental setup, we clarify that Minimality and Sufficiency share a joint estimation step, similar to Disentanglement and Completeness (DCI), which is why they are displayed as pairs. The results highlight three regimes: (i) Factor-VAEand SAP are the most efficient, maintaining negligible and nearly constant execution times; (ii) DCIsuffers from super-linear scaling costs; and (iii) our proposed metrics, while scaling almost linearly like MIG and Modularity, demonstrate superior efficiency, requiring less computation time than these linear baselines.
@@ 6 Neural Representations Experiments
To this end, we train models known to induce disentangled representations on three widely used datasets: DSprites (Matthey et al., 2017), MPI3D (Gondal et al., 2019), and Shapes3D (Burgess and Kim, 2018). For each dataset, we explore six encoder architectures (detailed in Appendix E) combined with six disentanglement losses: $\beta$-VAE (Higgins et al., 2017), AnnealedVAE (Burgess et al., 2018), $\beta$-TCVAE (Chen et al., 2018), FactorVAE (Kim and Mnih, 2018), Ada-GVAE (Locatello et al., 2020), and HFSVAE (Roth et al., 2023), yielding 36 encoders per dataset. All models are trained using the hyperparameters listed in Appendix D.
@@ 6.2 Interventional Robustness
Another motivation for disentangled representations is their robustness to interventions on the underlying factors. Specifically, if a factor $y_{i}$ is modified while all others remain fixed, the resulting change in the representation should be confined to its corresponding dimension $z_{j}$. This property is particularly critical for controlled generative models: if it holds, modifying $z_{j}$ ensures that the generated data changes exclusively in the associated factor $y_{i}$. To quantify this, Suter et al. (2019) introduced the Interventional Robustness Score (IRS). Formally, IRS evaluates the causal link between a factor $y_{i}$ and a latent dimension $z_{j}$ by comparing post-interventional distributions. A score of $1.0$ implies that $z_{j}$ is robust: its value changes if and only if we intervene on $y_{i}$ (i.e., $P(z_{j}|do(y_{i}))\neq P(z_{j}|do(y^{\prime}_{i}))$), but remains invariant if we intervene on any other factor $y_{k\neq i}$. We next compute the rank correlation between IRS and all considered metrics.
@@
Figure 6: Rank correlation between disentanglement metrics and Interventional Robustness Score across different levels of factor dependence. Legend as in Figure 5.
@@ 7 Conclusions
8 Limitations and Future Directions
While our framework provides a formal and rigorous definition of disentanglement, we acknowledge two primary limitations that open avenues for future research.
Differentiability and Optimization.
First, the proposed metrics rely on discrete binning to compute mutual information, rendering them non-differentiable. Consequently, they currently serve as evaluation tools rather than objective functions for training. A promising direction to address this is the use of variational approximations to derive differentiable upper bounds on the mutual information terms, as demonstrated in Alemi et al. (2017). Similarly, adopting an approach analogous to Almudévar et al. (2025b) could enable the direct optimization of models towards the minimality and sufficiency criteria.
Scalability to High-Dimensional Factors.
Second, the discretization-based approximation does not scale well to scenarios where the factors of variation or representations are high-dimensional. As the dimensionality increases, the number of bins required grows exponentially, making the estimation intractable. To extend our framework to complex, high-dimensional factors, future work could leverage more robust multivariate estimators, such as the Kraskov-Stögbauer-Grassberger estimator (Kraskov et al., 2004) or the Usable Information metric (Xu et al., 2020).
@@ Acknowledgments
This work has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101007666, MCIN/AEI/10.13039/501100011033 under Grant PID2024-155948OB-C53, and the Government of Aragón (Grant Group T36 23R).
Code Availability
An implementation accompanying this work is available at https://github.com/antonioalmudevar/dependent_disentanglement